# WebAccessVL: Making an Accessible Web via Violation-Conditioned VLM

### v1

@@ Abstract
We present a vision-language model (VLM) that automatically edits website HTML to address Web Content Accessibility Guidelines 2 (WCAG2) violations. We formulate this as a supervised image-conditioned program synthesis task, where the model learns to correct HTML given the HTML and its rendering. We collected WebAccessVL, a new dataset with manually corrected accessibility violations, establishing paired training data. We then propose a violation-conditioned VLM that additionally conditions on the WCAG2 violation count to guide the correction process. Experiments demonstrate that our method effectively reduces the average number of violations from 5.34 to 0.44 per website, outperforming commercial LLM APIs (Gemini, GPT-5). A perceptual study confirms that our edited websites maintain the original visual appearance and content.
@@ 1 Introduction
An accessible web is “perceivable, operable, understandable, and robust” to everyone, including people with disabilities as defined by WCAG2 [1]. Certain color combinations, such as red and green, may not be distinguished by people with protanopia (color blindness) and should be avoided. Web accessibility ensures that everyone has equal access to information and is legally required in the US (through the ADA [39]) or in the EU (through the EAA [12]). While required by law, surveys by WebAIM [43, 42] found that “95.9% of home pages had detected WCAG2 failures” and that the developers often “lacked the awareness and skills” to design WCAG2 compliant websites. We show that advances in VLMs can help tackle these challenges.
@@
There has been growing interest in leveraging artificial intelligence (AI) to improve web accessibility, such as employing large language models (LLMs) to automatically evaluate accessibility compliance [2]. Other efforts focus on correcting accessibility issues in websites and mobile interfaces by prompting LLMs [17, 30]. These approaches rely exclusively on text-based reasoning without incorporating any visual information. That is, they overlook the rendered appearance of a website, the primary way in which HTML code is perceived by users. Some accessibility violations are more apparent in the rendered view, and less so in the code. This observation motivates us to integrate website renderings as part of the input to the model.
@@
In this work, we formulate the task of making a website accessible as a computer vision problem of image-conditioned program synthesis. Given an HTML document and a screenshot of its rendering, the goal is to produce a modified HTML of the “same website” while ensuring WCAG2 compliance. See Fig. 1 for illustration. The main challenges include (a) designing a model that is aware of the WCAG2 guidelines and applies appropriate modifications; (b) maintaining the output HTML to be faithful to the input website design and rendering; (c) developing a suitable benchmark and metrics to quantify models’ performance.
@@
In this work, we create a dataset consisting of 2,500 webpage HTMLs, of which we manually correct each HTML to better follow WCAG2 standards. Each HTML takes around 7 to 10 minutes to edit. For evaluation, we introduce quantitative metrics to study the accessibility (# of violations) and faithfulness to the original input (similarity in rendering). Finally, we benchmark existing open-API and open-weight LLMs and VLMs.
With paired data, a baseline is to directly perform supervised fine-tuning (SFT), i.e., HTML as input, and generate a more WCAG2-compliant HTML. Differently, we propose to train the model to be violation aware. Unlike SFT, we introduce the number of violations as an additional input to the VLM. As the model is violation-aware, at test-time, we incorporate negative guidance sampling to encourage the model to generate HTML with fewer violations; See Fig. 2 for an overview. Finally, we conduct extensive experiments across 17 models, including LLMs and VLMs, and train our violation-conditioned VLM on open-source models to assess the effectiveness. Our proposed model reduces violations by 73.9% compared with GPT-5, which has the lowest number of violations out of the baselines.
@@
  • We propose a violation-conditioned VLM, which adapts pre-trained VLMs to improve web accessibility.
  • Through extensive experiments, we show that VLMs outperform LLM for this task and the effectiveness of our approach over baselines.
Figure 2: Overview of our image-conditioned program (HTML) synthesis pipeline. Our model takes in an HTML $\bm{x}$}, the rendered website image $\bm{I}$}, and the violation condition $\bm{c}$}. The Vision-Language Model processes the input tokens and vision tokens extracted from $\bm{I}$} to generate output tokens autoregressively. Our negative guidance sampling refines the token logits to reduce the number of violations in the output, ensuring better compliance with WCAG 2.2 accessibility guidelines.
@@ 2 Related Work
AI for accessibility. Existing work has leveraged AI to tackle accessibility challenges, e.g., automated tools designed to suggest source code fixes detected by checkers in mobile apps [30]. Ahmed et al. [2] evaluate ChatGPT’s ability to generate and improve web pages in line with WCAG and show that prompting only resolves simple issues, but struggles with complex tasks. Other works have also employed LLMs, such as ChatGPT, to detect [28] or fix violations through a prompting-based method [33, 3, 17, 5]. In contrast, we introduce a violation-conditioned model that is fine-tuned on a dedicated accessibility dataset created in this work. Importantly, we take both visual and textual input. This allows our method to generate HTML that follows the accessibility guidelines while preserving the original design, providing a more effective solution for web accessibility.
Visual programming and HTML synthesis. Early work by Andreas et al. [4] tackles the problem of visual question answering by assembling neural modules to build an end-to-end differentiable network, one of the first works to output a program from an input image. More recently, VisProg [14] proposes an approach that tackles complex, compositional visual tasks when given natural language instructions and outputs a program that may call other off-the-shelf computer vision models to perform the task, demonstrating impressive zero-shot reasoning capability. Differently, this work focuses on generating accessible HTML while preserving website design.
Research on HTML synthesis primarily focuses on converting rendered website images into code. Beltramelli [7] introduce an end-to-end method for generating code from graphical user interface screenshots using CNNs and RNNs. Building upon VLMs, Lee et al. [22] leverage a pre-trained VLM to transform webpage screenshots into simplified HTML. More recently, the synthetic WebSight dataset [21] provides HTML code paired with corresponding screenshots to facilitate research in this domain, and Roberts et al. [35] introduce a benchmark for evaluating VLMs on extracting structural information from images, including HTML code. Recent studies present datasets and benchmarks for HTML code synthesis using VLMs [49, 45], but these works do not focus on accessibility.
@@
Next, Qwen VL Series [40, 34] supports multiple languages and can handle multiple images concurrently during training, building on the Qwen language models [46]. Many other deep-net architectures have also been proposed for VLMs. Models such as OtterHD [23], mPLUG-Owl [48], and InternLM-XComposer2-4KHD [9] are designed to utilize high-resolution images, while LLaVA-NeXT [27], Mini-Gemini [24], and MM1 [29] divide the input image into multiple crops.
We show that prompting off-the-shelf VLMs yields some gains in accessibility compliance. In contrast, SFT on our proposed dataset, WebAccessVL, combined with the negative guidance, significantly improves the model’s ability to correct accessibility violations.
@@ 3 Approach
Our goal is to build a model that refines the HTML of a website to be WCAG2-compliant while maintaining the website’s content and design. To accomplish this goal, we collected a dataset for SFT (Sec. 3.1). We then propose a violation-conditioned VLM that adapts a pre-trained model for this task (Sec. 3.2).
@@ 3.1 WebAccessVL Dataset
Dataset construction methodology. To construct a set of websites’ HTML, we randomly sampled 2,500 websites from a large-scale HTML dataset provided by Laurençon et al. [21]. During sampling, we verify the availability of essential assets (e.g., images and icons). Websites with missing assets that led to rendering errors were excluded. To ensure reproducibility and prevent similar issues, we save all required assets locally.
@@
To annotate the HTML files, an annotator modifies the code and repeatedly renders each webpage to ensure that its visual design remains consistent with the original version. We use an industrial-grade accessibility checker developed by IBM [19] to minimize accessibility violations to the best of their ability. On average, this process takes approximately 7–10 minutes per webpage. Notably, due to the technical complexity of the task, the annotator requires an advanced degree in computer science.
@@
Dataset statistics. We analyzed the distribution of the collected websites, focusing on the types of accessibility violations and their dependence on visual versus textual information. Among the 26 violation categories, we identified 8 that correspond to vision-related tasks requiring visual understanding. Our analysis shows that 35.8% of violations involve vision-related factors, while the remaining 64.2% are purely language-based, indicating that visual information cannot be overlooked. Below, we briefly describe the major violation types; additional details are provided in Appendix Sec. A1.
Table 1: Accessibility violation statistics.
Violation Percentage (%)
Language violations
     aria_content_in_landmark $25.08$
     html_lang_exists $23.66$
     page_title_exists $23.66$
     skip_main_exists $5.94$
     text_block_heading $5.75$
Vision violations
     text_contrast_sufficient $76.63$
     img_alt_valid $15.90$
     style_color_misuse $6.49$
     caption_track_exists $0.51$
     svg_graphics_labelled $0.47$
@@
Vision violations. In Tab. 1, we present the five most frequent vision-related accessibility issues in the dataset, ranked by their occurrence rates: ① Text contrast sufficient violation means that the ratio of text contrast and background is less than a 3:1 contrast ratio. ② Img alt valid violation is triggered when alternative text is missing. ③ The checker flags a style color misuse violation when there are only color differences to mark required fields, i.e., users who cannot perceive the color will not be able to perform further actions. ④ caption track exists violation occurs when there is no textual information for video content. ⑤ SVG graphics labeled violation is similar to a missing alt text, but it differs in terms of format.
@@
Language violations. As shown in Tab. 1, the top five occurring language-related violations in the dataset are: ❶ aria content in landmark violation is when there is no role or subsection specified. ❷ html lang exists violation is triggered when no language is specified on the website. ❸ page title exists violation is marked if there is no head element and title element in HTML code. ❹ skip main exists violation happens when main tag is missing. ❺ text block heading violation is when there are no headings, i.e., the website has dense blocks of text.
3.2 Violation-conditioned VLM
Problem formulation. Given the input HTML of a website $\bm{x}$} and its rendering $\bm{I}$}, the model $p_{\theta}$ aims to output a revised HTML $\bm{y}$} of the same website but with fewer accessibility violations under WCAG2. To make the model aware of the violations, we introduce a binary indicator ${\bm{c}}\in\{c_{=0},c_{>0}\}$ that specifies whether the refined HTML $\bm{y}$} still violates any WCAG2 guidelines (i.e., whether the number of violations is greater than zero). Conceptually, the model is an image-conditioned program generator. The model is trained by minimizing the negative log-likelihood over the training set $\mathcal{D}$}, i.e.,
VLM details. Our violation-conditioned VLM is adapted from a pre-trained autoregressive token-based model, where textual and visual inputs are tokenized through text/vision encoders. The model then autoregressively generates a sequence of output tokens, which are subsequently decoded into text.
To prepare the inputs for the VLM, we prepend the input HTML $\bm{x}$} with the prompt “Modify the HTML code to comply with WCAG 2.2 Accessibility Guidelines” and encode it into text tokens ${\bm{t}}=({\bm{t}}_{1},{\bm{t}}_{2},\dots,{\bm{t}}_{L})$. Without the prepended prompt, we found that the models’ outputs are likely not in HTML format. Next, the rendered image $\bm{I}$} is processed by the vision encoder to obtain vision tokens ${\bm{v}}=({\bm{v}}_{1},\dots,{\bm{v}}_{K})$. The violation indicator $\bm{c}$} is tokenized using the text encoder, where $c_{=0}$ is represented with text prompt “The expected output HTML has zero violations” and $c_{>0}$ is represented with “The expected output HTML has non-zero violations”. The target HTML $\bm{y}$} is also encoded into text tokens ${\bm{u}}=({\bm{u}}_{1},{\bm{u}}_{2},\dots,{\bm{u}}_{T})$. SFT of the VLM $\tilde{p}_{\theta}$ is performed by minimizing the token-level negative log-likelihood:
where $\tilde{{\mathcal{D}}}$ denotes the tokenized training set.
Negative guidance sampling. To generate an HTML without violations $\bm{y}$}, we could set ${\bm{c}}=c_{=0}$, sample from $\tilde{p}_{\theta}$, and decode the set of generated tokens. In practice, we found that using a more advanced sampling strategy further reduces the number of violations. Specifically, we incorporate LLM classifier-free guidance with a negative prompt [36, 10, 15, 37]. Let
denote the conditional logit for the token ${\bm{u}}_{t}$. Negative guidance sampling subtracts a penalty term from the conditional logit, discouraging tokens associated with non-zero violation states.
Formally, the negative guided logit for the token ${\bm{u}}_{t}$ is
$\hskip-28.45274pt\ell^{\text{NG}}_{\theta}({\bm{u}}_{t},\gamma)\triangleq\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},c_{>0})$
$+\gamma\cdot\Bigl(\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},c_{=0})-\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},c_{>0})\Bigr),$ (4)
where $\gamma\in{\mathbb{R}}_{+}$ is the guidance scale. Intuitively, when $\gamma=1$, the sampling reduces to sampling from the conditional distribution with ${\bm{c}}=c_{=0}$. For $\gamma>1$, the logit of tokens associated with a violation greater than zero, i.e., $c_{>0}$, is being downweighted.

### v2

# WebAccessVL: Violation-Aware VLM for Web Accessibility

@@ Abstract
We present a vision-language model (VLM) that automatically edits website HTML to address violations of the Web Content Accessibility Guidelines 2 (WCAG2) while preserving the original design. We formulate this as a supervised image-conditioned program synthesis task, where the model learns to correct HTML given both the code and its visual rendering. We create WebAccessVL, a website dataset with manually corrected accessibility violations. We then propose a violation-conditioned VLM that further takes the detected violations’ descriptions from a checker as input. This conditioning enables an iterative checker-in-the-loop refinement strategy at test time. We conduct extensive evaluation on both open API and open-weight models. Empirically, our method achieves 0.211 violations per website, a 96.0% reduction from the 5.34 violations in raw data and 87% better than GPT-5. A perceptual study also confirms that our edited websites better maintain the original visual appearance and content.
@@ 1 Introduction
An accessible web is “perceivable, operable, understandable, and robust” to everyone, including people with disabilities as defined by WCAG2 [1]. Certain color combinations, such as red and green, may not be distinguished by people with protanopia (partial color blindness) and should be avoided. Web accessibility ensures that everyone has equal access to information and is legally required in the US (through the ADA [40]) or in the EU (through the EAA [15]). While required by law, surveys by WebAIM [44, 43] found that “95.9% of home pages had detected WCAG2 failures” and that the developers often “lacked the awareness and skills” to design WCAG2 compliant websites. We show that advances in VLMs can help tackle these challenges.
@@
There has been growing interest in leveraging artificial intelligence (AI) to improve web accessibility, for example, by employing large language models (LLMs) to automatically evaluate accessibility compliance [2]. Other efforts focus on correcting accessibility issues in websites and mobile interfaces by prompting LLMs [20, 31]. These approaches rely exclusively on text-based reasoning without incorporating any visual information. That is, they overlook the rendered appearance of a website, the primary way in which users perceive HTML code. Some accessibility violations are more apparent in the rendered view, and less so in the code. This observation motivates us to integrate website renderings as part of the input to the model.
@@
We formulate the task of making a website accessible as a computer vision problem of image-conditioned program synthesis. Given an HTML document and a screenshot of its rendering, the goal is to produce a modified HTML of the “same website” while ensuring WCAG2 compliance (see Fig. 1 for illustration). The main challenges include (a) designing a model that is aware of the WCAG2 guidelines and applies appropriate modifications; (b) maintaining the output HTML to be faithful to the input website design and rendering; (c) developing a suitable benchmark and metrics to quantify models’ performance.
@@
In this work, we create a dataset consisting of 1,500 webpage HTMLs. We then manually correct websites’ HTML to better follow WCAG 2.0 standards and use these as ground-truth annotations. Each HTML takes around 7 to 10 minutes to edit. For evaluation, we introduce quantitative metrics to study the accessibility (# of violations) and faithfulness to the original input (similarity in rendering). Finally, we benchmark existing open-API and open-weight LLMs and VLMs.
For paired data, a baseline is to perform supervised fine-tuning (SFT), i.e., take an HTML document as input and generate a more WCAG2-compliant HTML document. In contrast, we propose training the model to be violation-aware by conditioning on parsed violation reports that describe specific issues detected in the input HTML. This conditioning has two key benefits: (a) it guides the model to focus on editing the flagged elements rather than generic accessibility edits that may inadvertently edit unrelated components, and (b) it enables an iterative checker-in-the-loop refinement strategy (at test time) where the model receives feedback on the remaining violations and progressively corrects them. Additionally, we incorporate classifier-free guidance to amplify the effect of violation conditioning; See Fig. 2 for an overview. We conduct extensive experiments across 17 models, including LLMs and VLMs. Our violation-conditioned VLM achieves 0.211 violations per website (Gemma 3) and 0.244 (Llama 3.2 Vision), an 87% reduction compared with GPT-5 (1.68 violations), which has the lowest count among baselines. More importantly, we maintain 90% structural accuracy versus GPT-5’s 0.5%, demonstrating that our method preserves design while fixing accessibility.
@@
  2. 2.
    We propose a violation-conditioned VLM, which adapts pre-trained VLMs for web accessibility. The design incorporates the violation description into the conditioning and hence enables a test-time refinement approach to further improve performance.
  3. 3.
    Through extensive experiments, we show that VLMs outperform LLMs for this task and the effectiveness of our approach over baselines.
Figure 2: Overview of our violation-conditioned VLM pipeline. Input Preparation: The input HTML $\bm{x}$} is analyzed by an accessibility checker to generate a violation report $\bm{c}$}, which is encoded with $\bm{x}$} into input tokens $\bm{t}$}. VLM Forward Pass: The website rendering $\bm{I}$} is processed by a vision encoder to obtain vision tokens $\bm{v}$}. The vision and language model autoregressively generates output tokens $\bm{u}$} conditioned on $\bm{t}$} and $\bm{v}$}. Negative Guidance Sampling: We apply negative guidance sampling to refine the token logits, amplifying the effect of the violation condition $\bm{c}$} to reduce violations in the generated HTML $\bm{y}$}. Checker-in-the-Loop: The checker iteratively refines the output by feeding remaining violations back into the model for further correction.
@@ 2 Related Work
AI for accessibility. Traditional accessibility checking relies on rule-based systems [10, 16]. They require manual maintenance of specialized constraints per violation type, and cannot reason about visual-semantic context, e.g., generating descriptive alt-text from images. To address these challenges, existing work has leveraged AI to tackle accessibility challenges, e.g., automated tools designed to suggest source code fixes detected by checkers in mobile apps [31]. Ahmed et al. [2] evaluate ChatGPT’s ability to generate and improve web pages in line with WCAG and show that prompting only resolves simple issues, but struggles with complex tasks. Other works have also employed LLMs, such as ChatGPT, to detect [29] or fix violations through a prompting-based method [34, 3, 20, 5]. In contrast, we introduce a violation-conditioned model that is fine-tuned on a dedicated accessibility dataset created in this work. Importantly, we take both visual and textual input. This allows our method to generate HTML that follows the accessibility guidelines while preserving the original design, providing a more effective solution for web accessibility.
Visual programming and HTML synthesis. Early work by Andreas et al. [4] tackles the problem of visual question answering by assembling neural modules to build an end-to-end differentiable network, one of the first works to output a program from an input image. More recently, VisProg [17] proposes an approach that tackles complex, compositional visual tasks when given natural language instructions and outputs a program that may call other off-the-shelf computer vision models to perform the task, demonstrating impressive zero-shot reasoning capability. In contrast, this work focuses on generating accessible HTML while preserving the website design.
Research on HTML synthesis primarily focuses on converting rendered website images into code. Beltramelli [8] introduce an end-to-end method for generating code from graphical user interface screenshots using CNNs and RNNs. Building upon VLMs, Lee et al. [23] leverage a pre-trained VLM to transform webpage screenshots into simplified HTML. More recently, the synthetic WebSight dataset [22] provides HTML code paired with corresponding screenshots to facilitate research in this domain, and Roberts et al. [36] introduced a benchmark for evaluating VLMs on extracting structural information from images, including HTML code. Recent studies present datasets and benchmarks for HTML code synthesis using VLMs [51, 46], but these works do not focus on accessibility.
@@
Next, Qwen-VL Series [41, 35] supports multiple languages and can handle multiple images concurrently during training, building on the Qwen language models [48]. Many other deep-net architectures have also been proposed for VLMs. Models such as OtterHD [24], mPLUG-Owl [50], and InternLM-XComposer2-4KHD [11] are designed to utilize high-resolution images, while LLaVA-NeXT [27], Mini-Gemini [25], and MM1 [30] divide the input image into multiple crops. We show that prompting off-the-shelf VLMs yields some gains in accessibility compliance. In contrast, SFT on our proposed dataset, WebAccessVL, combined with the negative guidance, significantly improves the model’s ability to correct accessibility violations.
@@ 3 Approach
Our goal is to build a model that refines the HTML of a website to be WCAG2-compliant while maintaining the website’s content and design. To accomplish this, we propose a violation-conditioned VLM that adapts a pre-trained model for this task (Sec. 3.1). We then describe the dataset collected for SFT (Sec. 3.2).
3.1 Violation-conditioned VLM
Problem formulation. Given the input HTML of a website $\bm{x}$} and its rendering $\bm{I}$}, the model $p_{\theta}$ aims to output a revised HTML $\bm{y}$} of the same website but with fewer accessibility violations under WCAG2. To guide the repair process, we condition the model on a parsed violation report $\bm{c}$} that describes the specific WCAG2 violations detected in the input HTML $\bm{x}$}. This report is obtained by running an accessibility checker (e.g., IBM’s checker [21]) on $\bm{x}$}, which identifies violation types, affected elements, and their locations in the code. Conceptually, the model is an image-conditioned program generator. The model is trained by minimizing the negative log-likelihood over the training set $\mathcal{D}$}, i.e.,
VLM details. Our violation-conditioned VLM is adapted from a pre-trained autoregressive token-based model, in which textual and visual inputs are tokenized by text/vision encoders. The model then autoregressively generates a sequence of output tokens, which are subsequently decoded into text.
To prepare the inputs for the VLM, we construct a prompt that includes: (i) a structured violation report $\bm{c}$} listing the violations detected in the input HTML $\bm{x}$}, (ii) the instruction “The expected output HTML has zero violations”, and (iii) the input HTML $\bm{x}$} itself. The entire prompt is encoded into text tokens ${\bm{t}}=({\bm{t}}_{1},{\bm{t}}_{2},\dots,{\bm{t}}_{L})$. The violation report $\bm{c}$} is serialized into a structured text format describing each violation with its rule ID, message, and location (e.g., “Violation 1: text_contrast_sufficient, Issue: Text contrast is 2.1:1, Location: //div[@id=’header’]”). Without the prepended instruction, we found that the model’s outputs are likely not in HTML format. Next, the rendered image $\bm{I}$} is processed by the vision encoder to obtain vision tokens ${\bm{v}}=({\bm{v}}_{1},\dots,{\bm{v}}_{K})$. During training, $\bm{c}$} describes the violations detected in the input HTML $\bm{x}$}, and the model learns to generate the corrected HTML $\bm{y}$} that fixes these violations. The target HTML $\bm{y}$} is also encoded into text tokens ${\bm{u}}=({\bm{u}}_{1},{\bm{u}}_{2},\dots,{\bm{u}}_{T})$. SFT of the VLM $\tilde{p}_{\theta}$ is performed by minimizing the token-level negative log-likelihood:
where $\tilde{{\mathcal{D}}}$ denotes the tokenized training set.
Negative guidance sampling. To generate an HTML without violations $\bm{y}$}, we condition on the detected violation report $\bm{c}$} from the input HTML, sample from $\tilde{p}_{\theta}$, and decode the set of generated tokens. In practice, we found that using a more advanced sampling strategy further reduces the number of violations. Specifically, we incorporate LLM classifier-free guidance [37, 13, 18, 38]. Let
denote the conditional logit for the token ${\bm{u}}_{t}$. During training, we randomly drop the violation report with probability $p_{\text{uncond}}=0.1$ to enable classifier-free guidance. At inference, we use negative guidance sampling to amplify the effect of conditioning on violations.
Formally, the negative guided logit for the token ${\bm{u}}_{t}$ is
$\ell^{\text{NG}}_{\theta}({\bm{u}}_{t},\gamma)\triangleq\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{c}}_{\emptyset})+\gamma\cdot\Bigl(\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{c}})-\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{c}}_{\emptyset})\Bigr),$ (4)
where $\gamma\in{\mathbb{R}}_{+}$ is the guidance scale, $\bm{c}$} denotes the violation report from the input HTML, and $\bm{c}$$_{\emptyset}$ denotes an unconditional prompt that instructs the model to return the HTML as-is without modifications. Intuitively, when $\gamma=1$, the sampling reduces to sampling from the conditional distribution with $\bm{c}$}. For $\gamma>1$, the model amplifies the effect of conditioning on the specific violations, upweighting tokens that lead to corrected HTML and downweighting tokens associated with unchanged output. We explored alternative unconditional prompt formulations and empirically determined this configuration to be most effective (see Sec. A3).
$\tilde{p}^{\text{NG}}_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{t}},{\bm{v}},{\bm{c}})={\exp\Bigl(\ell^{\text{NG}}_{\theta}({\bm{u}}_{t},\gamma)\Bigr)}\Big/{\sum_{{\bm{u}}^{\prime}\in{\mathcal{V}}}\exp\Bigl(\ell^{\text{NG}}_{\theta}({\bm{u}}^{\prime},\gamma)\Bigr)}.$ (5)
Checker-in-the-loop inference. The violation conditioning mechanism described above enables an iterative refinement strategy at inference time. Let $\mathcal{C}(\cdot)$ denote the accessibility checker that maps HTML to its corresponding violation report. We define an iterative process where, at iteration $i$, we generate:
${\bm{y}}^{(i)}\sim p_{\theta}(\cdot\mid{\bm{x}}^{(i-1)},{\bm{I}}^{(i-1)},{\bm{c}}^{(i-1)}),\quad{\bm{c}}^{(i)}=\mathcal{C}({\bm{y}}^{(i)}),$ (6)
with initial conditions ${\bm{x}}^{(0)}={\bm{x}}$, ${\bm{I}}^{(0)}={\bm{I}}$, and ${\bm{c}}^{(0)}=\mathcal{C}({\bm{x}})$. This process continues for $i=1,\dots,K$ or until convergence, defined as either $|{\bm{c}}^{(i)}|=0$ or ${\bm{y}}^{(i)}={\bm{y}}^{(i-1)}$, where $|{\bm{c}}^{(i)}|$ denotes the number of violations.
Crucially, without violation conditioning, the checker’s output cannot be meaningfully fed back into the model; in this case, the model would simply receive the same HTML input repeatedly without guidance on what remains to be fixed. By conditioning on the specific violations detected at each iteration, the model receives targeted feedback on which accessibility issues persist and can better focus its corrections accordingly. In practice, we find that most samples converge in the second pass, with minimal improvement from additional passes.
@@ 3.2 WebAccessVL Dataset
Dataset construction methodology. To construct a set of websites’ HTML, we randomly sampled 1,500 websites from a large-scale HTML dataset provided by Laurençon et al. [22]. During sampling, we verify the availability of essential assets (e.g., images and icons). Websites with missing assets that led to rendering errors were excluded. To ensure reproducibility, we have saved all required assets locally.
@@
To annotate the HTML files, an annotator modifies the code and repeatedly renders each webpage to ensure that its visual design remains consistent with the original version. We use an industrial-grade accessibility checker developed by IBM [21] to minimize accessibility violations to the best of the annotator’s ability. On average, this process takes approximately 7 to 10 minutes per webpage. Notably, due to the technical complexity of the task, the annotator requires an advanced degree in computer science.
@@
Dataset statistics. We analyzed the distribution of the collected websites, focusing on the types of accessibility violations and their dependence on visual versus textual information. Among the 26 violation categories, we identified 8 that correspond to vision-related tasks requiring visual understanding. Our analysis shows that 35.8% of violations involve vision-related factors, while the remaining 64.2% are purely language-based, indicating that visual information should not be overlooked. Below, we briefly describe the major violation types; additional details are provided in Appendix Sec. A1.
Refer to caption Refer to caption Refer to caption
Figure 3: Pie charts showing the distribution of total violations, vision violations, and language violations. While the majority of the violations are language-based, more than a third of the total violations are vision-based. The most common vision violation is insufficient text contrast (“text contrast sufficient”), and the most common language violation is “aria content in landmark”.
@@
Vision violations. Fig. 3 shows the five most frequent vision-related accessibility issues in the dataset, ranked by occurrence rate: ① Text contrast sufficient violation means that the text-to-background contrast ratio is less than 3:1. ② Img alt valid violation is triggered when alternative text is missing. ③ The checker flags a style color misuse violation when there are only color differences to mark required fields, i.e., users who cannot perceive the color will not be able to perform further actions. ④ caption track exists violation occurs when there is no textual information for video content. ⑤ SVG graphics labeled violation is similar to a missing alt text, but it differs in terms of format.
@@
Language violations. As shown in Fig. 3, the top five occurring language-related violations in the dataset are: ❶ aria content in landmark violation is when there is no role or subsection specified. ❷ html lang exists violation is triggered when no language is specified on the website. ❸ page title exists violation is marked if there is no head element and title element in HTML code. ❹ skip main exists violation happens when main tag is missing. ❺ text block heading violation is when there are no headings, i.e., the website has dense blocks of text.
4 Experiments
We evaluate our violation-conditioned VLM on the WebAccessVL dataset. We benchmark open API and open weight models, then evaluate supervised fine-tuning on our dataset. Ablation studies validate our design choices, and a perceptual study checks whether quantitative metrics match human judgment.
4.1 Experimental Setup
[…]
Evaluation. We measure whether models improve accessibility while preserving the original design. We report four metrics:
  1. 1.
    Accessibility guidance violations $(\downarrow)$: Average number of WCAG2 violations detected by IBM’s accessibility checker [21].
  2. 2.
  3. 3.
  4. 4.
    Tree edit distance $(\downarrow)$: Tree edit distance between input and output HTML. Using the DOM structure, we treat each HTML tag (<div>, <body>, <nav>) as a tree node. Insert, update, and remove operations each cost one.
4.2 Main Results
We compare our approach against 17 baseline models across three categories: (1) commercial API models (GPT-4o [32], GPT-5 [33], Claude 3.5 Sonnet [6], Gemini 1.5 Flash [39]), (2) open source LLMs (Llama 3.2 [14], Qwen 2/2.5/3 [48, 49, 47]), and (3) open source VLMs (Llama 3.2 Vision [14], Qwen 2/2.5/3 VL [41, 35, 7], LLaVA 1.5 [26]/1.6 [27], DeepSeek VL 2 [45], Janus Pro [9]). All baselines are prompted with detailed instructions on WCAG2 guidelines and asked to modify the HTML to improve accessibility while maintaining visual consistency.
Tab. 1 presents the main results. The raw input websites average 5.335 violations, consistent with WebAIM [44] findings on widespread accessibility issues.
The accessibility preservation tradeoff. API models show a tradeoff between reducing violations and preserving design. GPT-5 reduced violations to 1.681 but rebuilds websites from scratch. Structural accuracy is near zero (0.005), and tree edit distance is high (15.573). GPT-5 appears to interpret the task as “generate an accessible website” rather than “fix this specific website.” GPT-4o Vision and Claude 3.5 maintain better structure (0.408 and 0.183 accuracy) but have more violations (2.918 and 2.310). Gemini 1.5 struggles on both.
Open-source VLMs fall behind APIs. Pre-trained open-source VLMs (without fine-tuning) perform poorly, retaining nearly all input violations (4 to 5 violations). Three reasons: (1) these models lack exposure to HTML rendering paired data during pre-training, (2) accessibility guidelines are specialized domain knowledge rarely seen in web corpora, and (3) the models struggle with long HTML documents. Even large recent VLMs like Qwen 3 VL (5.227 violations) barely improve over raw input. Text-only LLMs show comparable or worse performance than their VLM counterpart when used in the zero-shot manner (Qwen 3: 5.185 vs 5.227 violations). In the SFT setting, VLMs outperform the text-only LLMs; see Tab. 3.
[…]